URGENT PLEASE HELP!!! I need help with bA certain commodity cost 0.6 pound sterling in the United Kingdom. In U.S. denominated dollars, this same item can be purchas
Answer #1

Given

1 pound Sterling = $ 1.9 USD

Answer a)

Purchase in UK

Cost = 0.6 pound Sterling

In USD terms

1 pound Sterling = $ 1.9 USD

0.6 pound Sterling = $ 0.6 * 1.9

0.6 pound Sterling = 1.14 USD

---------------------------

Purchase in US

Cost = 80 Cents = $ 0.8 USD

In US we get the commodity at $ 0.8 USD while in UK the same we get in $ 1.14 USD.

Hence, it is cheaper is US and we should buy from US

-------------------------------------------------------------------------------

Answer b)

Quantity to be purchased = 100,000

Purchase in US

1 commodity = $ 0.8 USD

100,000 commodity = 100,000 * 0.8 = $ 80,000 USD

Purchase in UK

1 commodity = $ 1.14 USD (found in answer (a))

100,000 commodity = 100,000 * 1.14 = $ 114,000 USD

Hence, Savings : 114,000 - 80,000 = $ 34000 USD